Solve for x
x=(x-2)(square root of x)

Answer by stanbon(75887)

Solve for x
x=(x-2)(square root of x)
---
Divide both sides by sqrt(x) to get:
sqrt(x) = x-2)
---
Square both sides:
x = x^2-4x+4
----
Rearrange:
x^2-5x+4 = 0
Factor:
(x-4)(x-1) = 0
x = 1 or x = 4
-----------------------------
Cheers,
Stan H.